OpenStudy (anonymous):

helppp please<3 4. What is the measure of an exterior angle of a regular octagon? (1 point) a. 360 b. 135 c. 60 d. 45

OpenStudy (anonymous):

What is the value of x in the regular polygon below? a. 40 b.120 c. 60 d. 150

OpenStudy (anonymous):

2. What is the measure of one interior angle of a regular 12-gon? (1 point) a. 30 b. 180 c. 1800 d. 150* my answer

OpenStudy (anonymous):

1. What is the sum of the interior angle measures of a 20-gon? (1 point) a. 360 b. 3240 c. 3600 d. 162* my answer

OpenStudy (anonymous):

45 is the answer

OpenStudy (anonymous):

for which one

OpenStudy (yrelhan4):

The exterior angles of a polygon always add up to 360 degrees. If you have a regular octagon, it will have 8 sides (and obviously 8 exterior angles). 360 / 8 = 45° degrees.

OpenStudy (yrelhan4):

Each exterior angle of a regular12-sided polygon measures 360/12=30 degrees Therefore each interior angle measures 180-30=150 degrees

OpenStudy (yrelhan4):

sum of an interior angle of a regular polygon = 180º(n – 2) for 20-gon = 180 * (20 -2) = 3240º

OpenStudy (yrelhan4):

done?

OpenStudy (yrelhan4):

What is the value of x in the regular polygon below? A Regular Polygon is a polygon with all sides and all angles equal. so 360/6=30degrees each angle. so x=10 degrees.